# Self-consistent Hartree-Fock-Bogoliubov approach for bosons: self-eliminating divergence and pure pair condensate

### v1

@@ Abstract
We investigate the thermodynamic properties of an interacting Bose gas with a condensate within the energy-functional formulation of the Hartree-Fock-Bogoliubov approach. For a contact interaction, we derive a self-consistent solution to the HFB equations that intrinsically eliminates divergence. This solution characterizes the equilibrium state featuring a condensate of correlated pairs of particles. We analyze the temperature dependence of key thermodynamic quantities such as condensate density, chemical potential, entropy, pressure, specific heat capacity at constant volume, and isothermal compressibility and compare them with predictions from the Popov approximation. Our results reveal a first-order phase transition between the normal state and the degenerate state, accompanied by an increased transition temperature relative to the ideal Bose gas. Analysis of the compressibility indicates that a pure pair condensate is unstable, and the stable equilibrium corresponds to a mixture of single-particle and pair condensates.
I Introduction
@@
The phenomenon of Bose-Einstein condensation in an ideal gas has a long history, originating with the works of Bose Bose (1924) and Einstein Einstein (1925) during the early stages of the development of quantum mechanics. The theoretical treatment of this phenomenon, even in the regime of weak interatomic interaction, proved to be a significant challenge. This is due to the well-known inapplicability of the thermodynamic perturbative approach, whose series expansion gives rise to divergent terms at zero momentum. The problem was solved by Bogoliubov for a homogeneous Bose gas, who developed a special technique based on replacing the creation and annihilation operators with zero momentum in the Hamiltonian by $c$-numbers Bogoliubov (1947). The Bogoliubov prescription for the field operators was employed to formulate the Gross-Pitaevskii approach Gross (1961); Pitaevskii (1961), which became a powerful tool for investigating inhomogeneous condensates Pethick and Smith (2008); Pitaevskii and Stringari (2016). Although both approaches are simple and effective in describing key features of Bose-Einstein condensation, their validity is restricted to zero temperature and asymptotically weak interactions, whereas all experiments on Bose-Einstein condensates in dilute gases are necessarily conducted at finite temperatures Anderson et al. (1995); Davis et al. (1995); Bradley et al. (1995). In addition, a more thorough investigation of the ground state of a weakly interacting Bose gas within the Bogoliubov approach reveals the presence of correlated pairs of particles with opposite momentum Bogolubov and Bogolubov Jr (2010); Akhiezer and Peletminskii (1981), which are analogous to Cooper pairs in the BCS theory Girardeau and Arnowitt (1959); Valatin (1961); Nozières, P. and Saint James, D. (1982). Therefore, a central problem lies in developing theoretical frameworks that incorporate both interparticle interactions and finite-temperature effects on the condensate properties.
@@
For a complete description of Bose-Einstein condensates at finite temperatures, it is necessary to use advanced perturbative methods and self-consistent approximations Zagrebnov and Bru (2001); Proukakis and Jackson (2008); Andersen (2004). In particular, an extension of the Bogoliubov theory to finite temperature, which formally leads to the elementary excitation spectrum at zero temperature but with a temperature-dependent condensate density, was developed by Popov Popov (1983, 1987). The Popov approximation (PA), which essentially ignores pair correlations, can be artificially derived from the more general Hartree–Fock–Bogoliubov (HFB) approach. In turn, the self-consistent HFB approach, in its various formulations Krasil’nikov and Peletminskii (1993); Akhiezer et al. (1994); Griffin (1996); Zagrebnov and Bru (2001); Poluéktov (2002); Proukakis and Jackson (2008); Andersen (2004), includes not only normal (single-particle) averages but also anomalous averages that describe both the single-particle condensate and the condensate of correlated pairs of particles. These pairs affect the thermodynamic properties of the system, the stability of the condensate, and the structure of the excitation spectrum. However, both the PA and the HFB approach are known to suffer from divergences in physical quantities at high momenta (ultraviolet divergence), resulting from the assumption of a contact interaction. Generally, there are two ways to overcome the problem – to employ the model potential, whose Fourier transform $U({\mathbf{p}})$ goes to zero for ${\mathbf{p}}\to\infty$ Caballero-Benítez et al. (2013); Bulakhov et al. (2018) or to renormalize the coupling constant associated with the $s$-wave scattering length Pethick and Smith (2008); Pitaevskii and Stringari (2016). In recent years, considerable attention has been devoted to the study of Bose–Einstein condensates at finite temperatures, as thermal effects play a crucial role in understanding the full thermodynamic behavior of such systems including the equation of state and the transition temperature shift due to interparticle interaction Mordini et al. (2020); Ota and Giorgini (2020); Vianello and Salasnich (2024).
@@
The main purpose of this paper is to study finite-temperature effects in a weakly interacting Bose gas within a self-consistent, divergent-free HFB approach that includes pair anomalous averages. To this end, in Section 2, we present the energy-functional formulation of the HFB approach developed in Krasil’nikov and Peletminskii (1993); Akhiezer et al. (1994). Section 3 is devoted to deriving a system of coupled equations for the energy functional corresponding to the pairwise interaction Hamiltonian. In Section 4, for a contact interaction potential, we obtain a solution with a self-eliminating divergence (SSED), which allows us to avoid procedures such as cut-off of integrals, the use of model potentials or renormalization. This solution corresponds to a condensate of correlated pairs, commonly referred to as a pair condensate. We also discuss some physical implications of the derived solution, such as zero pressure at zero temperature and the interpretation of the divergent term as an effective attraction. Section 5 focuses on the investigation of the temperature dependence of key thermodynamic quantities, such as condensate density, chemical potential, pressure, entropy, specific heat capacity, and isothermal compressibility within SSED and PA. In particular, we suggest a first-order phase transition between the normal state and the state with broken U(1) symmetry. We also observe an increase in the transition temperature compared to that of an ideal gas. We find that the isothermal compressibility indicates the instability of a pure pair condensate (or SSED), allowing us to conclude that the stable equilibrium corresponds to a mixture of single-particle and pair condensates. Finally, Section 6, summarizes and discusses the results.
II Self-consistent formalism and basic equations
@@
[…]
A direct application of this Hamiltonian to study equilibrium and non-equilibrium properties of many-body systems leads to significant and often insurmountable difficulties, already in the case of a normal state. Therefore, some approximate formulations, such as the thermodynamic perturbative approach or mean-field theories, need to be developed for the Hamiltonian under consideration Akhiezer and Peletminskii (1981); Abrikosov et al. (1963). In turn, for systems with broken symmetries, the emergence of additional order parameters further complicates their description, even in the regime of weak interatomic interactions.
The quasiparticle approach has proven to be a powerful framework for studying interacting systems with spontaneously broken symmetries. It is precisely this non-perturbative approach that we use to formulate the basic equations describing a superfluid Bose system with two distinct types of condensates. Following Krasil’nikov and Peletminskii (1993); Akhiezer et al. (1994), we introduce the general form of the statistical operator of an ideal non-equilibrium Bose gas with broken U(1) symmetry:
@@
${\rm Sp\,\rho a_{\mathbf{p}}}=b_{\mathbf{p}},\quad{\rm Sp\,\rho a^{\dagger}_{\mathbf{p}}}=b^{*}_{\mathbf{p}},$ (4)
@@
to eliminate the linear terms associated with $C_{\mathbf{p}}$ in the statistical operator (3). The quantity $Z$ is determined from the normalization condition ${\rm Sp}\,\rho=1$. Thus, as a result of the unitary transformation (5), we have
${\rm Sp}\,\rho a^{{\dagger}}_{\mathbf{p}_{1}}a_{\mathbf{p}_{2}}=f_{\mathbf{p}_{2}\mathbf{p}_{1}}+b^{*}_{\mathbf{p}_{1}}b_{\mathbf{p}_{2}},$ (6)
${\rm Sp}\,\rho a_{\mathbf{p}_{1}}a_{\mathbf{p}_{2}}=g_{\mathbf{p}_{2}\mathbf{p}_{1}}+b_{\mathbf{p}_{1}}b_{\mathbf{p}_{2}},\quad{\rm Sp}\,\rho a^{\dagger}_{\mathbf{p}_{1}}a^{\dagger}_{\mathbf{p}_{2}}=g^{\dagger}_{\mathbf{p}_{2}\mathbf{p}_{1}}+b^{*}_{\mathbf{p}_{1}}b^{*}_{\mathbf{p}_{2}},$ (7)
@@
$f_{\mathbf{p}_{2}\mathbf{p}_{1}}={\rm Sp}\,\rho^{(0)}a^{\dagger}_{\mathbf{p}_{1}}a_{\mathbf{p}_{2}},\quad g_{\mathbf{p}_{2}\mathbf{p}_{1}}={\rm Sp}\,\rho^{(0)}a_{\mathbf{p}_{1}}a_{\mathbf{p}_{2}},\quad g^{\dagger}_{\mathbf{p}_{2}\mathbf{p}_{1}}={\rm Sp}\,\rho^{(0)}a^{\dagger}_{\mathbf{p}_{1}}a^{\dagger}_{\mathbf{p}_{2}}$ (8)
and $\rho^{(0)}=U\rho U^{\dagger}.$ Consequently, the non-equilibrium state of the system under consideration is specified by the condensate amplitudes $b_{\mathbf{p}}$ and $b^{*}_{\mathbf{p}}$, as well as by the normal $f_{\mathbf{p}_{2}\mathbf{p}_{1}}$ ($f=f^{\dagger}$) and the anomalous $g_{\mathbf{p}_{2}\mathbf{p}_{1}}$ and $g^{\dagger}_{\mathbf{p}_{2}\mathbf{p}_{1}}$ ($g=\tilde{g}$) distribution functions, where the tilde symbol denotes transposition. The condensate amplitudes describe a single-particle (Bogoliubov) condensate, while the presence of pair anomalous averages may signal the formation of a condensate of correlated pairs. It should be noted that the distinction between a condensate of correlated pairs and a molecular condensate lies in the nature of the interparticle binding. In molecules, particles form tightly two-body bound states, which then condense as single objects Timmermans et al. (1999); Radzihovsky et al. (2008); Peletminskii et al. (2014, 2017). In contrast, correlated pairs, arising from interparticle interactions and quantum fluctuations in equilibrium, are neither localized nor long-lived in the sense of conventional molecules. Moreover, for a system with $N$ particles, the number of diatomic molecules is $N/2$, while the number of correlated pairs is $N(N-1)/2$. Therefore, the correlated pairs under consideration are similar to Cooper pairs in BCS theory without well-defined bound energy. In Fermi systems, a smooth transition from a condensate of molecules to a condensate of delocolized Cooper pairs, occurring as the attraction between particles is varied, is known as the BEC-BCS crossover Bloch et al. (2008); Zwerger (2011); Chen et al. (2024). For further analysis, it is convenient to combine the normal and anomalous distribution functions into a block matrix and the condensate amplitudes into a column vector:
@@
According to (8), the matrices $A$, $B$, and $B^{\dagger}$ that enter the statistical operator $\rho^{(0)}$ can be expressed in terms of the quantities $f$, $g$, and $g^{\dagger}$. Therefore, $\rho^{(0)}=\rho^{(0)}(\hat{f})$ and, consequently, the non-equilibrium entropy $S=-{\rm Sp}\,\rho^{(0)}\ln\rho^{(0)}$ is also a certain functional of the normal and anomalous distribution functions, but not of the condensate amplitudes, $S=S(\hat{f})$. Note that $\hat{f}$ is isomorphic under a canonical transformation, which simultaneously reduces both $\rho^{(0)}$ and $\hat{f}$ to a diagonal form. As a result, one can show that (for details, see Ref. Krasil’nikov and Peletminskii (1993); Akhiezer et al. (1994)):
@@
$E(\hat{f},\hat{b})={\rm Sp}\,\rho H,\quad N(f,\hat{b})={\rm Sp}\,\rho N,$ (11)
where $N(f,\hat{b})$ is independent of anomalous distribution functions. Then, the variational principle states that the equilibrium values of $\hat{f}$ and $\hat{b}$ should be found from the requirement of maximum entropy at fixed system energy and particle number. Therefore, introducing the Lagrange multipliers $\beta$ and $\mu$, we arrive at the following functional to be minimized:
$\mathcal{L}(\hat{f},\hat{b})=-S(\hat{f})+\beta E(\hat{f},\hat{b})-\mu N(f,\hat{b}).$ (12)
The solution of the formulated variational problem gives the following self-consistency equations: Krasil’nikov and Peletminskii (1993); Akhiezer et al. (1994):
@@
The derived equations determine the equilibrium values of the normal and anomalous distribution functions, as well as the condensate amplitude, for a given energy functional. Due to the bosonic structure of the operator $\hat{f}$ (see Eq. (13)), the Lagrange multipliers $\mu$ and $\beta=1/T$ should be interpreted as the chemical potential and the inverse temperature, respectively. Equation (14) is similar to the stationary Gross-Pitaevskii equation without external potential. However, the key distinction lies in the fact that the system’s energy now depends not only on the condensate amplitudes but also on the normal and anomalous distribution functions. It is worth noting that the quasiparticle formulation presented here is essentially equivalent to the HFB approach for bosons Andersen (2004); Poluéktov (2002). Moreover, it is constructed in close analogy to the Fermi liquid theory extended to superfluid states Krasil’nikov et al. (1990).
III Energy functional and coupled equations for homogeneous state

### v2

@@ Abstract
We investigate the thermodynamic properties of an interacting Bose gas with a condensate within the energy-functional formulation of the Hartree-Fock-Bogoliubov (HFB) approach. For a contact interaction, we derive a self-consistent solution to the HFB equations that intrinsically eliminates divergence. This solution characterizes the equilibrium state featuring a condensate of correlated pairs of particles. We analyze the temperature dependence of key thermodynamic quantities such as condensate density, chemical potential, entropy, pressure, specific heat capacity at constant volume, and isothermal compressibility and compare them with predictions from the Popov approximation (PA). We predict that the transition temperature shifts to higher values due interactions, with the HFB approach yielding a larger shift than the PA. Analysis of the compressibility indicates that a pure pair condensate is unstable, and the stable equilibrium corresponds to only a mixture of single-particle and pair condensates.
1 Introduction
@@
The phenomenon of Bose-Einstein condensation in an ideal gas has a long history, originating with the works of Bose Bose [1924] and Einstein Einstein [1925] during the early stages of the development of quantum mechanics. The theoretical treatment of this phenomenon, even in the regime of weak interatomic interaction, proved to be a significant challenge. This is due to the well-known inapplicability of the thermodynamic perturbative approach, whose series expansion gives rise to divergent terms at zero momentum. The problem was solved by Bogoliubov for a homogeneous Bose gas, who developed a special technique based on replacing the creation and annihilation operators with zero momentum in the Hamiltonian by $c$-numbers Bogoliubov [1947]. The Bogoliubov prescription for the field operators was employed to formulate the Gross-Pitaevskii approach Gross [1961], Pitaevskii [1961], which became a powerful tool for investigating inhomogeneous condensates Pethick and Smith [2008], Pitaevskii and Stringari [2016]. Although both approaches are simple and effective in describing key features of Bose-Einstein condensation, their validity is restricted to zero temperature and asymptotically weak interactions, whereas all experiments on Bose-Einstein condensates in dilute gases are necessarily conducted at finite temperatures Anderson et al. [1995], Davis et al. [1995], Bradley et al. [1995]. In addition, a more thorough investigation of the ground state of a weakly interacting Bose gas within the Bogoliubov approach reveals the presence of correlated pairs of particles with opposite momentum Bogolubov and Bogolubov Jr [2010], Akhiezer and Peletminskii [1981], which are analogous to Cooper pairs in the BCS theory Girardeau and Arnowitt [1959], Valatin [1961], Nozières, P. and Saint James, D. [1982]. Therefore, a central problem lies in developing theoretical frameworks that incorporate both interparticle interactions and finite-temperature effects on the condensate properties.
@@
For a complete description of Bose-Einstein condensates at finite temperatures, it is necessary to use advanced perturbative methods and self-consistent approximations Shi and Griffin [1998], Morgan [2000], Zagrebnov and Bru [2001], Proukakis and Jackson [2008], Andersen [2004]. In particular, an extension of the Bogoliubov theory to finite temperature, which formally leads to the elementary excitation spectrum at zero temperature but with a temperature-dependent condensate density, was developed by Popov Popov [1983, 1987]. The Popov approximation (PA), which essentially ignores pair correlations, can be artificially derived from the more general Hartree–Fock–Bogoliubov (HFB) approach. In turn, the self-consistent HFB approach, in its various formulations Krasil’nikov and Peletminskii [1993], Akhiezer et al. [1994], Griffin [1996], Zagrebnov and Bru [2001], Poluéktov [2002], Proukakis and Jackson [2008], Andersen [2004], includes not only normal (single-particle) averages but also anomalous averages that describe both the single-particle condensate and the condensate of correlated pairs of particles. These pairs affect the thermodynamic properties of the system, the stability of the condensate, and the structure of the excitation spectrum. However, both the PA and the HFB approach are known to suffer from divergences in physical quantities at high momenta (ultraviolet divergence), resulting from the assumption of a contact interaction. Generally, there are two ways to overcome the problem – to employ the model potential, whose Fourier transform $U({\mathbf{p}})$ goes to zero for $\mathbf{p}$$\to\infty$ Caballero-Benítez et al. [2013], Bulakhov et al. [2018] or to renormalize the coupling constant associated with the $s$-wave scattering length Pethick and Smith [2008], Pitaevskii and Stringari [2016]. In recent years, considerable attention has been devoted to the study of Bose–Einstein condensates at finite temperatures, as thermal effects play a crucial role in understanding the full thermodynamic behavior of such systems including the equation of state and the transition temperature shift due to interparticle interaction Mordini et al. [2020], Ota and Giorgini [2020], Vianello and Salasnich [2024].
@@
The main purpose of this paper is to study finite-temperature effects in a weakly interacting Bose gas within a self-consistent, divergent-free HFB approach that includes pair anomalous averages. To this end, in Section 2, we present the energy-functional formulation of the HFB approach developed in Krasil’nikov and Peletminskii [1993], Akhiezer et al. [1994]. Section 3 is devoted to deriving a system of coupled equations for the energy functional corresponding to the pairwise interaction Hamiltonian. In Section 4, for a contact interaction potential, we obtain a solution with a self-eliminating divergence (SSED), which allows us to avoid procedures such as cut-off of integrals, the use of model potentials or renormalization. This solution corresponds to a condensate of correlated pairs, commonly referred to as a pair condensate. We also discuss some physical implications of the derived solution, such as zero pressure at zero temperature and the interpretation of the divergent term as an effective attraction. Section 5 focuses on the investigation of the temperature dependence of key thermodynamic quantities, such as condensate density, chemical potential, pressure, entropy, specific heat capacity, and isothermal compressibility within SSED and PA. In particular, we suggest a first-order phase transition between the normal state and the state with broken U(1) symmetry. We also observe an increase in the transition temperature compared to that of an ideal gas. We find that the isothermal compressibility indicates the instability of a pure pair condensate (or SSED), allowing us to conclude that the stable equilibrium corresponds to a mixture of single-particle and pair condensates. Finally, Section 6, summarizes and discusses the results.
2 Self-consistent formalism and basic equations
@@
$U({\mathbf{p}})=\int{d\mathbf{x}}V(\mathbf{x})e^{-i\mathbf{px}/\hbar}.$ (2)
A direct application of this Hamiltonian to study equilibrium and non-equilibrium properties of many-body systems leads to significant and often insurmountable difficulties, already in the case of a normal state. Therefore, some approximate formulations, such as the thermodynamic perturbative approach or mean-field theories, need to be developed for the Hamiltonian under consideration Akhiezer and Peletminskii [1981], Abrikosov et al. [1963]. In turn, for systems with broken symmetries, the emergence of additional order parameters further complicates their description, even in the regime of weak interatomic interactions.
The quasiparticle approach has proven to be a powerful framework for studying interacting systems with spontaneously broken symmetries. It is precisely this non-perturbative approach that we use to formulate the basic equations describing a superfluid Bose system with two distinct types of condensates. Following Krasil’nikov and Peletminskii [1993], Akhiezer et al. [1994], we introduce the general form of the statistical operator of an ideal non-equilibrium Bose gas with broken U(1) symmetry:
@@
${\rm Tr\,\rho a_{\mathbf{p}}}=b_{\mathbf{p}},\quad{\rm Tr\,\rho a^{\dagger}_{\mathbf{p}}}=b^{*}_{\mathbf{p}},$ (4)
@@
to eliminate the linear terms associated with $C_{\mathbf{p}}$ in the statistical operator (3). The quantity $Z$ is determined from the normalization condition $\textrm{Tr}\,\rho=1$. Thus, as a result of the unitary transformation (5), we have
$\textrm{Tr}\,\rho a^{{\dagger}}_{\mathbf{p}_{1}}a_{\mathbf{p}_{2}}=f_{\mathbf{p}_{2}\mathbf{p}_{1}}+b^{*}_{\mathbf{p}_{1}}b_{\mathbf{p}_{2}},$ (6)
$\textrm{Tr}\,\rho a_{\mathbf{p}_{1}}a_{\mathbf{p}_{2}}=g_{\mathbf{p}_{2}\mathbf{p}_{1}}+b_{\mathbf{p}_{1}}b_{\mathbf{p}_{2}},\quad\textrm{Tr}\,\rho a^{\dagger}_{\mathbf{p}_{1}}a^{\dagger}_{\mathbf{p}_{2}}=g^{\dagger}_{\mathbf{p}_{2}\mathbf{p}_{1}}+b^{*}_{\mathbf{p}_{1}}b^{*}_{\mathbf{p}_{2}},$ (7)
@@
$f_{\mathbf{p}_{2}\mathbf{p}_{1}}=\textrm{Tr}\,\rho^{(0)}a^{\dagger}_{\mathbf{p}_{1}}a_{\mathbf{p}_{2}},\quad g_{\mathbf{p}_{2}\mathbf{p}_{1}}=\textrm{Tr}\,\rho^{(0)}a_{\mathbf{p}_{1}}a_{\mathbf{p}_{2}},\quad g^{\dagger}_{\mathbf{p}_{2}\mathbf{p}_{1}}=\textrm{Tr}\,\rho^{(0)}a^{\dagger}_{\mathbf{p}_{1}}a^{\dagger}_{\mathbf{p}_{2}}$ (8)
and $\rho^{(0)}=U\rho U^{\dagger}.$ Consequently, the non-equilibrium state of the system under consideration is specified by the condensate amplitudes $b_{\mathbf{p}}$ and $b^{*}_{\mathbf{p}}$, as well as by the normal $f_{\mathbf{p}_{2}\mathbf{p}_{1}}$ ($f=f^{\dagger}$) and the anomalous $g_{\mathbf{p}_{2}\mathbf{p}_{1}}$ and $g^{\dagger}_{\mathbf{p}_{2}\mathbf{p}_{1}}$ ($g=\tilde{g}$) distribution functions, where the tilde symbol denotes transposition. The condensate amplitudes describe a single-particle (Bogoliubov) condensate, while the presence of pair anomalous averages may signal the formation of a condensate of correlated pairs. It should be noted that the distinction between a condensate of correlated pairs and a molecular condensate lies in the nature of the interparticle binding. In molecules, particles form tightly two-body bound states, which then condense as single objects Timmermans et al. [1999], Radzihovsky et al. [2008], Peletminskii et al. [2014, 2017]. In contrast, correlated pairs, arising from interparticle interactions and quantum fluctuations in equilibrium, are neither localized nor long-lived in the sense of conventional molecules. Moreover, for a system with $N$ particles, the number of diatomic molecules is $N/2$, while the number of correlated pairs is N(N1)/2N(N-1)/2. Therefore, the correlated pairs under consideration are similar to Cooper pairs in BCS theory without well-defined bound energy. In Fermi systems, a smooth transition from a condensate of molecules to a condensate of delocolized Cooper pairs, occurring as the attraction between particles is varied, is known as the BEC-BCS crossover Bloch et al. [2008], Zwerger [2011], Chen et al. [2024]. For further analysis, it is convenient to combine the normal and anomalous distribution functions into a block matrix and the condensate amplitudes into a column vector:
@@
[…]
@@
$E(\hat{f},\hat{b})=\textrm{Tr}\,\rho H,\quad N(f,\hat{b})=\textrm{Tr}\,\rho N,$ (11)
[…]
The solution of the formulated variational problem gives the following self-consistency equations: Krasil’nikov and Peletminskii [1993], Akhiezer et al. [1994]:
@@
The derived equations determine the equilibrium values of the normal and anomalous distribution functions, as well as the condensate amplitude, for a given energy functional. Due to the bosonic structure of the operator $\hat{f}$ (see Eq. (13)), the Lagrange multipliers $\mu$ and $\beta=1/T$ should be interpreted as the chemical potential and the inverse temperature, respectively. Equation (14) is similar to the stationary Gross-Pitaevskii equation without external potential. However, the key distinction lies in the fact that the system’s energy now depends not only on the condensate amplitudes but also on the normal and anomalous distribution functions. It is worth noting that the quasiparticle formulation presented here is essentially equivalent to the HFB approach for bosons Andersen [2004], Poluéktov [2002]. Moreover, it is constructed in close analogy to the Fermi liquid theory extended to superfluid states Krasil’nikov et al. [1990].
3 Energy functional and coupled equations for homogeneous state